In a survey of 100 out patients who were reported at a hospital one day found out that 70 complained of fever,50 complained of stomach ache and 30 were injured. All 100 patients had at least one of the complaints and 44 had exactly two of the three complaints. How many patients had all three complaints.

Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

F-70
S-50
I=30
two complaints=44=F+S, S+I, F+I
150 complaints altogether.
88 of them are with 44 patients
That leaves 56 patients.
x have 3 complaints
56-x have 1 complaint
The total complaints for that group is 3x+56-x, and that equals 62 complaints
2x+56=62
2x=6
x=3 patients with all 3 complaints ANSWER
3 have 3 for total of 9
44 have 2 for total of 88
53 have 1 for total of 53

Answer by ikleyn(52803)

.
Let F be the set and the number of those who complained of fever        (F = 70);
    S be the set and the number of those who complained of stomach ache (S = 50);
    I be the set and the number of those who wee injured                (I = 30).


Do not worry that I denoted by the same symbol the set and the number: I made it for simplicity, 
and you always can distinct from the context what I am talking about.


Let FS be the intersection of the sets F and S and the number of elements in this set at the same time.

Let FI be the intersection of the sets F and I and the number of elements in this set at the same time.

Let SI be the intersection of the sets S and I and the number of elements in this set at the same time.



    Let FSI = x be the intersection of the sets F, S and I and the number of elements in this intersection at the same time.

    I called the last quantity as "x", since it is our major unknown in this problem.


SO,        the set FS includes those who has two complains F and S, but also includes those who has all 3 complains. (*)
Similarly, the set FI includes those who has two complains F and I, but also includes those who has all 3 complains. (*)
Finally,   the set SI includes those who has two complains S and I, but also includes those who has all 3 complains. (*)

   
From the elementary theory of finite sets, we have this equation 

100 = F + S + I - FS - FI - SI + x.       (1)


From the given part of the condition, we have this equation

(FS-x) + (FI-x) + (SI-x) = 44.            (2)


    Indeed,    (FS-x) are those and only those who has exactly two complains F and S and do not have third complain.  (*)
    Similarly, (FI-x) are those and only those who has exactly two complains F and I and do not have third complain.  (*)
    Finally,   (SI-x) are those and only those who has exactly two complains S and I and do not have third complain.  (*)
    So, the left part of (2) are those and only those who have exactly two complains - exactly as its right side.     (*)


Now, we can re-write equation (1) in this form

100 = 70 + 50 + 30 - (FS-x) - (FI-x) - (SI-x) -3x + x = 

    = 150 - [(FS-x) + (FI-x) + (SI-x)] - 2x = 150 - 44 - 2x = 106 - 2x,

which gives us 

2x = 106 - 100 = 6.


Hence,  x = 3.


Answer.  The number of patients who had all three complaints was 3.

For the key equation (1) see my lessons
    - Counting elements in sub-sets of a given finite set
and especially
    - Advanced problems on counting elements in sub-sets of a given finite set
in this site.

-------------------

This problem was posted to the forum long time ago, and I solved it under this link

https://www.algebra.com/algebra/homework/sets-and-operations/sets-and-operations.faq.question.1096428.html

https://www.algebra.com/algebra/homework/sets-and-operations/sets-and-operations.faq.question.1096428.html


This time I placed additional explanations marked (*) in this post, to make the solution more clear.
The major idea of the solution remained unchangeable.